3/x-2 + 5/x+1 = 1/x^2-x-2
I think this is considered fractional equations and inequalities?

Answer by stanbon(75887)

3/x-2 + 5/x+1 = 1/x^2-x-2
-----------
The least common multiple of the denominators is x^2-x-2
Multiply thru by the lcm to get:
3(x+1) + 5(x-2) = 1
8x-7 = 1
8x = 8
x = 1
=================
Cheers,
Stan H.